A jet airliner moving initially at 399 mph (with respect to the ground) to the east moves into a region where the wind is blowing at 552 mph in a direction 72◦ north of east. What is the new speed of the aircraft with respect to the ground? Answer in units of mph.


1
Expert's answer
2020-10-02T07:14:04-0400

Solution

Given angle "\theta=72^0"

Speed of aircraft

"V=\sqrt{v_a^2+v_w^2+2v_av_w\cos\theta}"

Where "v_a, v_w" are speed of airlines and wind with respect to ground.

Therefore new speed of airlines ith respect to ground can be expressed as

"V=\sqrt{(399)^2+(552)^2+2\times (399)\times (552)cos72^°}"

"V=774.6 mph"